Solve the method of undermined coefficient (D^2+8D+16)y=8e^-2x ; y(0)= 2 y'(0)= 0


1
Expert's answer
2021-10-18T16:39:17-0400

Homogeneous differential equation


"(D^2+8D+16)y=0"

Corresponding (auxiliary) equation


"r^2+8r+16=0"

"(r+4)^2=0"

"r_1=r_2=-4"

The general solution of the homogeneous differential equation is


"y_h=c_1xe^{-4x}+c_2e^{-4x}"

Find the particular solution of the nonhomogeneous differential equation


"y_p=Ae^{-2x}"

"y_p'=-2Ae^{-2x}"

"y_p''=4Ae^{-2x}"

Substitute


"4Ae^{-2x}-16Ae^{-2x}+16Ae^{-2x}=8e^{-2x}"

"A=2"

The general solution of the nonhomogeneous differential equation is


"y=c_1xe^{-4x}+c_2e^{-4x}+2e^{-2x}"

"y(0)= 2, y'(0)= 0"


"c_1(0)e^{-4(0)}+c_2e^{-4(0)}+2e^{-2(0)}=2=>c_2=0"

"y=c_1xe^{-4x}+2e^{-2x}"

"y'=c_1e^{-4x}-4c_1xe^{-4x}-4e^{-2x}"

"c_1e^{-4(0)}-4c_1(0)e^{-4(0)}-4e^{-2(0)}=0=>c_1=4"

The solution of the given initial value problem is


"y=4xe^{-4x}+2e^{-2x}"